# Stability of ranks under field extensions

### v3

@@ 5. Partition rank and slice rank under field extension
Next we consider the slice rank of a linear subspace of tensors. Given a linear subspace $\mathsf{W}\subseteq\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$ over a field $\mathbb{K}$, the slice rank of $\mathsf{W}$ is defined as
[…]
@@
$\mathsf{V}_{i}\coloneqq\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{i-1}}\otimes\mathbb{K}^{n_{i+1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}},$ (23)
@@
Let $\mathbb{K}$ be a field. For a linear subspace $\mathsf{W}\subseteq\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$, $\operatorname{SR}_{\mathbb{K}}(\mathsf{W})$ is the minimal positive integer $r$ to ensure the existence of $r_{i}\in\mathbb{N}$ and $a_{i1},\dots,a_{ir_{i}}\in\mathbb{K}^{n_{i}}$ for $1\leq i\leq d$, such that $r=\sum_{i=1}^{d}r_{i}$ and every $T\in\mathsf{W}$ admits a decomposition
@@
We denote by $s$ the minimal integer such that (24) holds. We prove that $s=t\coloneqq\operatorname{SR}_{\mathbb{K}}(\mathsf{W})$. If $\langle\mathsf{W},\mathsf{U}_{1}\otimes\cdots\otimes\mathsf{U}_{d}\rangle=0$, then $\mathsf{W}\subseteq\sum_{i=1}^{d}\mathsf{U}_{i}^{\perp}\otimes\mathsf{V}_{i}$ where $\mathsf{U}_{i}^{\perp}\coloneqq((\mathbb{K}^{n_{i}})^{\ast}/\mathsf{U}_{i})^{\ast}$. Hence any $T\in\mathsf{W}$ admits a decomposition (24) if we choose $r_{i}=\dim\mathsf{U}_{i}^{\perp}$ and let $\{a_{i1},\dots,a_{ir_{i}}\}$ be a basis of $\mathsf{U}_{i}^{\perp}$ for $1\leq i\leq d$. Thus we have $t=\sum_{i=1}^{d}\dim\mathsf{U}_{i}^{\perp}=\sum_{i=1}^{d}r_{i}\geq s$.
@@
Then clearly we have $\langle\mathsf{W},\mathsf{U}_{1}\otimes\cdots\otimes\mathsf{U}_{d}\rangle=0$ and this implies $s=\sum_{i=1}^{d}r_{i}\geq t$. ∎
@@
For any field $\mathbb{K}$ and linear subspace $\mathsf{W}\subseteq\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$, there exist $n_{d+1},n_{d+2}\in\mathbb{N}$ and $T_{\mathsf{W}}\in\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d+2}}$ such that $\operatorname{SR}_{\mathbb{K}}(\mathsf{W})=\operatorname{SR}_{\mathbb{K}}(T_{\mathsf{W}})$. .
@@
[…]
$T_{\mathsf{W}}\coloneqq\sum_{j=1}^{n}T_{j}\otimes\left(\sum_{i=1}^{m}e_{ij}\otimes e_{ij}\right)\in\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}\otimes\mathbb{K}^{mn}\otimes\mathbb{K}^{mn}.$
@@
Let $T_{\mathsf{W}}=\sum_{i=1}^{d+2}\sum_{j=1}^{r_{i}}a_{ij}\otimes S_{ij}$ be a slice rank decomposition of $T_{\mathsf{W}}$, where $a_{ij}\in\mathbb{K}^{n_{i}}$ and $S_{ij}\in\mathsf{V}_{i}$ for $21\leq i\leq d+$ 2 and $1\leq j\leq r_{i}$. Here $\mathsf{V}_{i}$ is defined as in (23). We denote
$\mathsf{W}^{\prime}\coloneqq\sum_{i=1}^{d}\sum_{j=1}^{r_{i}}\left\langle a_{ij}\otimes S_{ij},(\mathbb{K}^{mn})^{\ast}\otimes(\mathbb{K}^{mn})^{\ast}\right\rangle\subseteq\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}.$
We observe that each $L\in\mathsf{W}^{\prime}$ is a linear combination of $a_{ij}\otimes S^{\prime}_{ij}$ for some
[…]
Without loss of generality, we may assume $T_{1},\cdots,T_{l}\in W^{\prime}$ and $T_{l+1},\cdots,T_{n}\notin W^{\prime}$ for some $0\leq l\leq n$. Therefore we have
$\operatorname{SR}_{\mathbb{K}}(\mathsf{W})\leq\sum_{i=1}^{d}r_{i}+\sum_{s=l+1}^{n}\max\{n_{i}\}_{i=1}^{d}\leq\sum_{i=1}^{d}r_{i}+(n-l)m.$
In the following, we regard $a_{ij}$’s (resp. $T_{k}$’s) as linear (resp. multilinear) polynomials. Since $T_{l+1},\dots,T_{n}\not\in\mathsf{W}^{\prime}$, we conclude that $T_{l+1}\cdots T_{n}\notin J$, where $J$ is the prime ideal generated by linear forms $a_{ij},1\leq i\leq d,1\leq j\leq r_{i}$. By Hilbert Nullstellensatz, there exist $v_{i}\in\overline{\mathbb{K}}^{n_{i}}$, $1\leq i\leq d$, such that
[…]
@@
$T_{W}(v_{1},\cdots,v_{d})$ $=\sum_{i=1}^{d+2}\sum_{j=1}^{r_{i}}(a_{ij}\otimes S_{ij})(v_{1},\dots,v_{d})$
@@
$T_{W}(v_{1},\cdots,v_{d})=\sum_{s=1}^{n}T_{s}(v_{1},\dots,v_{d})\left(\sum_{i=1}^{m}e_{is}\otimes e_{is}\right)=\sum_{s=l+1}^{n}T_{s}(v_{1},\dots,v_{d})\left(\sum_{i=1}^{m}e_{is}\otimes e_{is}\right).$ (26)
Both (25) and (26) are decompositions of the matrix $T_{W}(v_{1},\cdots,v_{d})\in\overline{\mathbb{K}}^{mn\times mn}$ into the sum of rank one matrices. Moreover, (26) is clearly a rank decomposition. This implies that $(n-l)m\leq r_{d+1}+r_{d+2}$ and
$\operatorname{SR}_{\mathbb{K}}(\mathsf{W})\leq\sum_{i=1}^{d}r_{i}+(n-l)m\leq\sum_{i=1}^{d+2}r_{i}=\operatorname{SR}_{\mathbb{K}}(T_{\mathsf{W}}).\qed$
@@
For any field $\mathbb{K}$ and linear subspaces $\mathsf{W}_{1}\subseteq\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$, $\mathsf{W}_{2}\subseteq\mathbb{K}^{m_{1}}\otimes\cdots\otimes\mathbb{K}^{m_{d}}$, we have
@@
Let $\mathbb{K}$ be a field. For any positive integer $d$ and linear subspace $\mathsf{W}\subseteq\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$, we have
@@
According to the proof of Lemma 5.8, there is a tensor $T_{\mathsf{W}}\in\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d+2}}$ such that $\operatorname{SR}_{\mathbb{K}}(\mathsf{W})=\operatorname{SR}_{\mathbb{K}}(T_{\mathsf{W}})$ and $\operatorname{SR}_{\overline{\mathbb{K}}}(\mathsf{W}\otimes\overline{\mathbb{K}})=\operatorname{SR}_{\overline{\mathbb{K}}}\left(T_{\mathsf{W}}^{\overline{\mathbb{K}}}\right)$. Now the desired inequality follows immediately from (20). ∎
We remark that Theorem 5.10 is proved for $d=2$ in [1] by a different approach. To conclude this section, we define for each $\mathsf{W}\subseteq\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$ and positive integer $1\leq k\leq\dim\mathsf{W}$ the following quantity:
@@
$\operatorname{SR}_{k,\mathbb{K}}(\mathsf{W})=\operatorname{SR}_{\mathbb{K}}(\mathsf{V})\leq 2\operatorname{SR}_{\overline{\mathbb{K}}}(\mathsf{V}\otimes\overline{\mathbb{K}})\leq 2\operatorname{SR}_{k,\overline{\mathbb{K}}}(\mathsf{W}\otimes\overline{\mathbb{K}}).\qed$
@@
$T_{s}\coloneqq I_{n}\otimes e_{k}\in\mathbb{K}^{n}\otimes\mathbb{K}^{n}\otimes\mathbb{K}^{n},\quad 1\leq s\leq n,$
@@
$T_{s}\in\mathsf{U}_{1}\otimes\mathbb{K}^{n}\otimes\mathbb{K}^{n}+\mathbb{K}^{n}\otimes\mathsf{U}_{2}\otimes\mathbb{K}^{n}+\mathbb{K}^{n}\otimes\mathbb{K}^{n}\otimes\mathsf{U}_{3},\quad 1\leq s\leq n,$
@@
$I_{n}=\langle T_{k_{0}},f\rangle\in\mathsf{U}_{1}\otimes\mathbb{K}^{n}+\mathbb{K}^{n}\otimes\mathsf{U}_{2}.$
@@
There is a constant $C\coloneqq C(d)>0$ such that for any linear subspace $\mathsf{W}\subseteq\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$ and positive integer $k\leq\mathsf{W}$, we have
$\operatorname{SR}_{k,\overline{\mathbb{K}}}(W\otimes\overline{\mathbb{K}})\leq C\operatorname{SR}_{k,\mathbb{K}}(\mathsf{W}).$
Acknowledgement
We would like to thank Daniel Zhu and Guy Moshkovitz for their helpful comments on the previous version of this manuscript. With their suggestions, we fill gaps in the proofs of Lemma 3.5 and Theorem 3.7.
@@ Appendix A Algebraic function field
We provide a brief introduction to the theory of algebraic function fields. Interested readers are referred to [40] for more details. An algebraic function field of one variable over $\mathbb{K}$ is a field extension $\mathbb{F}/\mathbb{K}$ such that there is a transcendental element $f\in\mathbb{F}$ and $[\mathbb{F}:\mathbb{K}(f)]<\infty$. We recall that any valuation ring $(\mathcal{O},\mathfrak{m},v_{\mathfrak{m}})$ such that $\mathbb{K}\subsetneq\mathcal{O}\subsetneq\mathbb{F}$ and $F(\mathcal{O})=\mathbb{F}$ is a discrete valuation ring (DVR). Here $\mathfrak{m}$ is the maximal ideal of $\mathcal{O}$, $F(\mathcal{O})$ is the field of fractions of $\mathcal{O}$ and $v_{\mathfrak{m}}$ is the valuation of $F(\mathcal{O})$. The maximal ideal $\mathfrak{m}$ of $\mathcal{O}$ is called a place of $\mathbb{F}/\mathbb{K}$. The degree of $\mathfrak{m}$ is defined as $\deg(\mathfrak{m})\coloneqq\dim_{\mathbb{K}}(\mathcal{O}/\mathfrak{m})$.
@@
[40, Theorem 1.5.17] Let $g$ be the genus of $\mathbb{F}/\mathbb{K}$. For each $D\in\operatorname{Div}(\mathbb{F}/\mathbb{K})$ with $\deg(D)\geq 2g-1$, we have $\ell(D)=\deg(D)+1-g$.
@@
[40, Theorem 1.4.11] All principal divisors have degree 00.
@@
[40, Theorem 1.3.1]] Let $\mathbb{F}/\mathbb{K}$ be a function field. Given distinct places $\mathfrak{m}_{1},\dots,\mathfrak{m}_{n}$ of $\mathbb{F}/\mathbb{K}$, $f_{1},\cdots,f_{n}\in\mathbb{F}$ and $r_{1},\dots,r_{n}\in\mathbb{Z}$, there exists some $f\in\mathbb{F}$ such that $v_{\mathfrak{m}_{i}}(f-f_{i})=r_{i}$ for each $1\leq i\leq n$.
@@
[40, Corollary 1.4.12] For any $D\in\operatorname{Div}(\mathbb{F}/\mathbb{K})$ with $\deg(D)<0$, we have $\ell(D)=0$.
@@
[2, Lemma 2.1] Let $q$ be a prime power. Let $\mathbb{F}/\mathbb{F}_{q}$ be an algebraic function field of genus $g$ containing at least $g+1$ places of degree one. Denote by $P_{1}(F/\mathbb{F}_{q})$ the set of places of degree one. For any $S\subseteq P_{1}(\mathbb{F}/\mathbb{F}_{q})$ with $|S|\geq g+1$, there exists a non-special divisor $D$ such that $\deg(D)=g-1$ and $supp(D)\subseteq S$. Here $supp(D)$ is the set of place with non-zero coefficient in $D$.
@@
[17, Theorem] For each positive integer $k$, there exists an algebraic function field $\mathbb{F}/\mathbb{F}_{q^{2}}$ whose genus is
@@
[40, Corollary 5.2.10] Any algebraic function field $\mathbb{F}/\mathbb{F}_{q}$ of genus $g$ admits at least one place of degree $r$ for each $r\geq 4g+3$.

### v4

@@ 5 Partition rank and slice rank under field extension
Next we consider the slice rank of a linear subspace of tensors. Given a linear subspace $\mathsf{W}\subseteq\mathbb{K}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d}}$ over a field $\mathbb{K}$, the slice rank of $\mathsf{W}$ is defined as
$\operatorname{SR}_{\mathbb{K}}(\mathsf{W})\coloneqq\min\left\{\sum_{i=1}^{d}\operatorname{codim}(\mathsf{U}_{i}):\langle\mathsf{W},\mathsf{U}_{1}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathsf{U}_{d}\rangle=0,\;\mathsf{U}_{i}\subseteq(\mathbb{K}^{n_{i}})^{\ast},\;1\leq i\leq d\right\}.$ (24)
@@
$\mathsf{V}_{i}\coloneqq\mathbb{K}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{i-1}}\otimes_{\mathbb{K}}\mathbb{K}^{n_{i+1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d}},$ (25)
@@
Let $\mathbb{K}$ be a field. For a linear subspace $\mathsf{W}\subseteq\mathbb{K}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d}}$, $\operatorname{SR}_{\mathbb{K}}(\mathsf{W})$ is the minimal positive integer $r$ to ensure the existence of $r_{i}\in\mathbb{N}$ and $a_{i1},\dots,a_{ir_{i}}\in\mathbb{K}^{n_{i}}$ for $1\leq i\leq d$, such that $r=\sum_{i=1}^{d}r_{i}$ and every $T\in\mathsf{W}$ admits a decomposition
@@
We denote by $s$ the minimal integer such that (26) holds. We prove that $s=t\coloneqq\operatorname{SR}_{\mathbb{K}}(\mathsf{W})$. If $\langle\mathsf{W},\mathsf{U}_{1}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathsf{U}_{d}\rangle=0$, then $\mathsf{W}\subseteq\sum_{i=1}^{d}\mathsf{U}_{i}^{\perp}\otimes\mathsf{V}_{i}$ where $\mathsf{U}_{i}^{\perp}\coloneqq((\mathbb{K}^{n_{i}})^{\ast}/\mathsf{U}_{i})^{\ast}$. Hence any $T\in\mathsf{W}$ admits a decomposition (26) if we choose $r_{i}=\dim\mathsf{U}_{i}^{\perp}$ and let $\{a_{i1},\dots,a_{ir_{i}}\}$ be a basis of $\mathsf{U}_{i}^{\perp}$ for $1\leq i\leq d$. Thus we have $t=\sum_{i=1}^{d}\dim\mathsf{U}_{i}^{\perp}=\sum_{i=1}^{d}r_{i}\geq s$.
@@
Then clearly we have $\langle\mathsf{W},\mathsf{U}_{1}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathsf{U}_{d}\rangle=0$ and this implies $s=\sum_{i=1}^{d}r_{i}\geq t$. ∎
@@
For any field $\mathbb{K}$ and linear subspace $\mathsf{W}\subseteq\mathbb{K}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d}}$, there exist $n_{d+1},n_{d+2}\in\mathbb{N}$ and $T_{\mathsf{W}}\in\mathbb{K}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d+2}}$ such that $\operatorname{SR}_{\mathbb{K}}(\mathsf{W})=\operatorname{SR}_{\mathbb{K}}(T_{\mathsf{W}})$.
@@
Assume $\dim\mathsf{W}=n$. Let $T_{1},\dots,T_{n}$ be a basis of $\mathsf{W}$. We denote $m\coloneqq\max\left\{n_{i}:1\leq i\leq d\right\}+1$. Suppose that $\{e_{ij}:1\leq i\leq m,\;1\leq j\leq n\}$ is a basis of $\mathbb{K}^{mn}$. We consider
$T_{\mathsf{W}}\coloneqq\sum_{j=1}^{n}T_{j}\otimes\left(\sum_{i=1}^{m}e_{ij}\otimes e_{ij}\right)\in\mathbb{K}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d}}\otimes_{\mathbb{K}}\mathbb{K}^{mn}\otimes_{\mathbb{K}}\mathbb{K}^{mn}.$
@@
Let $T_{\mathsf{W}}=\sum_{i=1}^{d+2}\sum_{j=1}^{r_{i}}a_{ij}\otimes S_{ij}$ be a slice rank decomposition of $T_{\mathsf{W}}$, where $a_{ij}\in\mathbb{K}^{n_{i}}$ and $S_{ij}\in\mathsf{V}_{i}$ for $1\leq i\leq d+2$ and $1\leq j\leq r_{i}$. Here $\mathsf{V}_{i}$ is defined as in (25). We denote
$\mathsf{W}^{\prime}\coloneqq\sum_{i=1}^{d}\sum_{j=1}^{r_{i}}\left\langle a_{ij}\otimes\mathsf{V}_{i},(\mathbb{K}^{mn})^{\ast}\otimes_{\mathbb{K}}(\mathbb{K}^{mn})^{\ast}\right\rangle\subseteq\mathbb{K}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d}}.$
We observe that each element of $\mathsf{W}^{\prime}$ is a linear combination of $a_{ij}\otimes R_{ij}$ for some
$R_{ij}\in\langle\mathsf{V}_{i},(\mathbb{K}^{mn})^{\ast}\otimes_{\mathbb{K}}(\mathbb{K}^{mn})^{\ast}\rangle,\;1\leq i\leq d,\;1\leq j\leq r_{i}.$
Without loss of generality, we may assume $T_{1},\cdots,T_{l}\in\mathsf{W}^{\prime}$ and $T_{l+1},\cdots,T_{n}\notin\mathsf{W}^{\prime}$ for some $0\leq l\leq n$. In the following, we regard $a_{ij}$’s (resp. $T_{k}$’s) as linear (resp. multilinear) polynomials. Since $T_{l+1},\dots,T_{n}\not\in\mathsf{W}^{\prime}$, we conclude that $T_{l+1}\cdots T_{n}\notin J$, where $J$ is the prime ideal generated by linear forms $a_{ij},1\leq i\leq d,1\leq j\leq r_{i}$. By Hilbert Nullstellensatz, there exist $v_{i}\in\overline{\mathbb{K}}^{n_{i}}$, $1\leq i\leq d$, such that
$a_{ij}(v_{i})=0,\;T_{s}(v_{1},\cdots,v_{d})\not=0,\;1\leq j\leq r_{i},\;l+1\leq s\leq d.$
@@
$T_{\mathsf{W}}(v_{1},\cdots,v_{d})$ $=\sum_{i=1}^{d+2}\sum_{j=1}^{r_{i}}(a_{ij}\otimes S_{ij})(v_{1},\dots,v_{d})$
@@
$T_{\mathsf{W}}(v_{1},\cdots,v_{d})=\sum_{s=1}^{n}T_{s}(v_{1},\dots,v_{d})\left(\sum_{i=1}^{m}e_{is}\otimes e_{is}\right)=\sum_{s=l+1}^{n}T_{s}(v_{1},\dots,v_{d})\left(\sum_{i=1}^{m}e_{is}\otimes e_{is}\right).$ (28)
Both (27) and (28) are decompositions of the matrix $T_{\mathsf{W}}(v_{1},\cdots,v_{d})\in\overline{\mathbb{K}}^{mn\times mn}$ into the sum of rank one matrices. Moreover, (28) is clearly a rank decomposition. This implies that $(n-l)m\leq r_{d+1}+r_{d+2}$ which implies $l=n$ since $r_{d+1}+r_{d+2}\leq\operatorname{SR}_{\mathbb{K}}(T_{\mathsf{W}})<m$ by construction. Hence all $T_{i}$ are in $\mathsf{W}^{\prime}$, and this implies $\operatorname{SR}_{\mathbb{K}}(\mathsf{W})\leq\operatorname{SR}_{\mathbb{K}}(\mathsf{W}^{\prime})\leq\operatorname{SR}_{\mathbb{K}}(T_{\mathsf{W}})$. ∎
@@
For any field $\mathbb{K}$ and linear subspaces $\mathsf{W}_{1}\subseteq\mathbb{K}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d}}$, $\mathsf{W}_{2}\subseteq\mathbb{K}^{m_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{m_{d}}$, we have
@@
Let $\mathbb{K}$ be a field. For any positive integer $d$ and linear subspace $\mathsf{W}\subseteq\mathbb{K}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d}}$, we have
@@
According to the proof of Lemma 5.9, there is a tensor $T_{\mathsf{W}}\in\mathbb{K}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d+2}}$ such that $\operatorname{SR}_{\mathbb{K}}(\mathsf{W})=\operatorname{SR}_{\mathbb{K}}(T_{\mathsf{W}})$ and $\operatorname{SR}_{\overline{\mathbb{K}}}(\mathsf{W}\otimes\overline{\mathbb{K}})=\operatorname{SR}_{\overline{\mathbb{K}}}\left(T_{\mathsf{W}}^{\overline{\mathbb{K}}}\right)$. Now the desired inequality follows immediately from (22). ∎
We remark that Theorem 5.11 is proved for $d=2$ in [1] by a different approach. To conclude this section, we define for each $\mathsf{W}\subseteq\mathbb{K}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d}}$ and positive integer $1\leq k\leq\dim\mathsf{W}$ the following quantity:
@@
$\operatorname{SR}_{k,\mathbb{K}}(\mathsf{W})=\operatorname{SR}_{\mathbb{K}}(\mathsf{V})\leq 2\operatorname{SR}_{\overline{\mathbb{K}}}(\mathsf{V}\otimes\overline{\mathbb{K}})\leq 2\operatorname{SR}_{k,\overline{\mathbb{K}}}(\mathsf{W}\otimes\overline{\mathbb{K}}).$
@@
$T_{s}\coloneqq I_{n}\otimes e_{k}\in\mathbb{K}^{n}\otimes_{\mathbb{K}}\mathbb{K}^{n}\otimes_{\mathbb{K}}\mathbb{K}^{n},\quad 1\leq s\leq n,$
@@
$T_{s}\in\mathsf{U}_{1}\otimes_{\mathbb{K}}\mathbb{K}^{n}\otimes_{\mathbb{K}}\mathbb{K}^{n}+\mathbb{K}^{n}\otimes_{\mathbb{K}}\mathsf{U}_{2}\otimes_{\mathbb{K}}\mathbb{K}^{n}+\mathbb{K}^{n}\otimes_{\mathbb{K}}\mathbb{K}^{n}\otimes_{\mathbb{K}}\mathsf{U}_{3},\quad 1\leq s\leq n,$
@@
$I_{n}=\langle T_{k_{0}},f\rangle\in\mathsf{U}_{1}\otimes_{\mathbb{K}}\mathbb{K}^{n}+\mathbb{K}^{n}\otimes_{\mathbb{K}}\mathsf{U}_{2}.$
@@
There is a constant $C\coloneqq C(d)>0$ such that for any linear subspace $\mathsf{W}\subseteq\mathbb{K}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d}}$ and positive integer $k\leq\mathsf{W}$, we have
$\operatorname{SR}_{k,\overline{\mathbb{K}}}(\mathsf{W}\otimes_{\mathbb{K}}\overline{\mathbb{K}})\leq C\operatorname{SR}_{k,\mathbb{K}}(\mathsf{W}).$
Appendix
@@ Appendix A Algebraic function field
We provide a brief introduction to the theory of algebraic function fields. Interested readers are referred to [46] for more details. An algebraic function field of one variable over $\mathbb{K}$ is a field extension $\mathbb{F}/\mathbb{K}$ such that there is a transcendental element $f\in\mathbb{F}$ and $\mathbb{F}:\mathbb{K}(f)]<\infty$. We recall that any valuation ring $(\mathcal{O},\mathfrak{m},v_{\mathfrak{m}})$ such that $\mathbb{K}\subsetneq\mathcal{O}\subsetneq\mathbb{F}$ and $F(\mathcal{O})=\mathbb{F}$ is a discrete valuation ring (DVR). Here $\mathfrak{m}$ is the maximal ideal of $\mathcal{O}$, $F(\mathcal{O})$ is the field of fractions of $\mathcal{O}$ and $v_{\mathfrak{m}}$ is the valuation of $F(\mathcal{O})$. The maximal ideal $\mathfrak{m}$ of $\mathcal{O}$ is called a place of $\mathbb{F}/\mathbb{K}$. The degree of $\mathfrak{m}$ is defined as $\deg(\mathfrak{m})\coloneqq\dim_{\mathbb{K}}(\mathcal{O}/\mathfrak{m})$.
@@
[46, Theorem 1.5.17] Let $g$ be the genus of $\mathbb{F}/\mathbb{K}$. For each $D\in\operatorname{Div}(\mathbb{F}/\mathbb{K})$ with $\deg(D)\geq 2g-1$, we have $\ell(D)=\deg(D)+1-g$.
@@
[46, Theorem 1.4.11] All principal divisors have degree 0.
@@
[46, Theorem 1.3.1] Let $\mathbb{F}/\mathbb{K}$ be a function field. Given distinct places $\mathfrak{m}_{1},\dots,\mathfrak{m}_{n}$ of $\mathbb{F}/\mathbb{K}$, $f_{1},\cdots,f_{n}\in\mathbb{F}$ and $r_{1},\dots,r_{n}\in\mathbb{Z}$, there exists some $f\in\mathbb{F}$ such that $v_{\mathfrak{m}_{i}}(f-f_{i})=r_{i}$ for each $1\leq i\leq n$.
@@
[46, Corollary 1.4.12] For any $D\in\operatorname{Div}(\mathbb{F}/\mathbb{K})$ with $\deg(D)<0$, we have $\ell(D)=0$.
@@
[4, Lemma 2.1] Let $q$ be a prime power. Let $\mathbb{F}/\mathbb{F}_{q}$ be an algebraic function field of genus $g$ containing at least $g+1$ places of degree one. Denote by $P_{1}(F/\mathbb{F}_{q})$ the set of places of degree one. For any $S\subseteq P_{1}(\mathbb{F}/\mathbb{F}_{q})$ with $|S|\geq g+1$, there exists a non-special divisor $D$ such that $\deg(D)=g-1$ and $supp(D)\subseteq S$. Here $supp(D)$ is the set of place with non-zero coefficient in $D$.
@@
[21, Theorem] For each positive integer $k$, there exists an algebraic function field $\mathbb{F}/\mathbb{F}_{q^{2}}$ whose genus is
@@
[46, Corollary 5.2.10] Any algebraic function field $\mathbb{F}/\mathbb{F}_{q}$ of genus $g$ admits at least one place of degree $r$ for each $r\geq 4g+3$.